The dollar fell from ¥106.85 to ¥102.65 over a one day period. Meanwhile, the Tokyo stock market fell 1.63% as measured in yen on the same day. What was the one-day dollar return on the Tokyo stock market?
Answer #1

On that day, the dollar value of the yen rose by (106.85 - 102.65)/102.65 = 4.09%.

R$ = [(1 + R¥) * (1 + Increase in dollar value of yen)] - 1

= [(1 - 0.0163)(1.0409)] - 1 = 1.0239 - 1 = 0.0239, or 2.39%